don willams uses his small motorboat to go 4 miles upstream to his favorite fishing spot. against the current, the trip takes 2/5 hour. with the current the trip takes 1/5 hour. how fast can the boat travel in still water? what is the speed of the current?
Answer by lwsshak3(11628)

don willams uses his small motorboat to go 4 miles upstream to his favorite fishing spot. against the current, the trip takes 2/5 hour. with the current the trip takes 1/5 hour. how fast can the boat travel in still water? what is the speed of the current?
***
let x=speed of boat in still water
let c=speed of current
x+c=speed of boat in downstream
x-c=speed of boat in upstream
distance= travel time*speed
(2/5)(x-c)=4
(1/5)(x+c)=4
.4x-.4c=4
.2x+.2c=4
..
.4x-.4c=4
.4x+.4c=8
add
.8x=12
x=15
.2c=4-.2x=4-3=1
c=5
speed of boat in still water=15 mph
speed of current=5 mph